A rocket of mass m = 20kg has M =180kg fuel. The uniform exhaust velocity of fuel is v(r)=1.6km/s (¡) calculate the minimum rate of consumption of fuel so that the rocket may rise from ground
 

Grade:11

1 Answers

Susmita
5 years ago
Force on the rocket is gravitational pull (m+M)g.
F=dP/dt=d{(m+M)v}/dt
m is constant and v has a fixed value in this problem.
F=vdm/dt
Or,dm/dt=F/v=(m+M)g/v=200×9.8/(1.6×1000)=1.225 kg/s
If it helps please approve.